# Improving Ab-Initio Cryo-EM Reconstruction with Semi-Amortized Pose Inference

### v1

@@ Appendix B Specialization of Encoder Heads
Figure 6: Average rotation error visualized over the unit sphere for different heads of our multi-head pose encoder ($M=4$). The sphere is uniformly divided into cells using HEALPix [53] and based on their ground-truth view-direction, images are assigned to corresponding cells. For each cell, the average rotation error is visualized, showing diverse behavior of different heads across the space. Blue and red colors show low and high error regions, respectively. Error ranges from zero to 180 degrees.
Appendix C More on Semi-Amortized vs. Fully-Amortized
@@
Appendix D More on Multi-Modal Pose Posterior
@@
Appendix E Videos
In the supplementary package, using ChimeraX [51] we provide videos that show reconstructions obtained by semi-amortized method, cryoAI and cryoSPARC on all synthetic and experimental datasets. In these 3D visualizations, we rotate the structure to show the resolved structure from different views.
Figure 7: Using four examples per dataset, the behavior of fully-amortized and semi-amortized pose inference methods are compared. Two rows correspond to Spliceosome and Spike datasets, respectively. Each plot shows the approximate log pose posterior, marginalized over in-plane rotations represented as a heat map on a uniform grid over the unit sphere $S^{2}$. Gnomonic projection to 2D is also applied, followed by zooming on the proximity of the mode of interest. Black dot is the starting point while blue dots and red dots show poses estimated by fully- and semi-amortized methods, respectively.

### v2

@@ Appendix C Specialization of Encoder Heads
Figure 6: Average rotation error is visualized over the sphere for different heads of the multi-head pose encoder ($M=4$). The sphere is uniformly divided into cells using HEALPix [53]; images are assigned to cells based on ground-truth view-direction. For each cell, the average rotation error is visualized, showing diverse behavior of different heads across the space. Blue and red colors show low and high error regions, respectively. Error ranges from zero to 180 degrees.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: We compare the behavior of fully-amortized and semi-amortized pose inference on four examples per dataset. Two rows correspond to Spike and Spliceosome datasets, respectively. Each plot shows the approximate log pose posterior, marginalized over in-plane rotations represented as a heat map on a uniform grid over the unit sphere $S^{2}$. Gnomonic projection to 2D is also applied, followed by zooming on the proximity of the mode of interest. Black cross is the starting point while blue dots and red dots show poses estimated by fully- and semi-amortized methods, respectively.
Appendix D More on Semi-Amortized vs. Fully-Amortized
@@
Appendix E More on Multi-Modal Pose Posterior